OpenStudy (anonymous):

Kirsten used a log to make an umbrella stand. The log was 28 in. tall and had a radius of 7.5 in. Kirsten cut a cylindrical shape, with a radius of 5.5 in., out of the center of the log. After Kirsten cuts out the cylinder, what is the volume of the remaining portion of the log? Use 3.14 to approximate pi, and express your answer to two decimal places.

OpenStudy (anonymous):

Find the volumes of the two cylinders.

OpenStudy (anonymous):

\(V = \pi r^2 h\) \(V = (3.14)(7.5^2)(28)\)

OpenStudy (anonymous):

Then find the volume of the other one: \(V = (3.14)(5.5^2)(28)\)